# Two-dimensional Entanglement-assisted Quantum Quasi-cyclic Low-density Parity-check Codes

### v1

@@ I Introduction
Let $H_{2\text{-}D}$ be the parity check tensor constructed based on the shifts in (8), and let $\ell_{1}$ and $\ell_{2}$ be any two horizontal layers in $H_{2\text{-}D}$, then we obtain
@@
Viewing the layers of $H_{2\text{-}D}$ as matrices, motivates the following definitions.
@@
The maximum number of linearly independent layers in a parity-check tensor is referred to as its rank.
Using this definition, we now state the rank of a parity-check tensor $H_{2\text{-}D}$ formed by selecting a subset of horizontal block-layers.
@@
Taking the shifts defined in (8) into account, it is straightforward to verify that the sum of all layers within any horizontal block-layer is the all-ones matrix, which implies a linear dependence among the layers in horizontal block-layers. Now, consider the $w$ horizontal block-layers of $\widetilde{H}_{2\text{-}D}$, and remove one horizontal layer from every horizontal block-layer except the first one. Now, consider the following linear combination of all the remaining layers in the all horizontal-block layers:
@@
where $L_{i}^{(j)}$ stands for the $i$th layer in the $j$th horizontal block-layer and the coefficients are binary scalars. Consider the inner product defined in (11) with $\pmod{2}$. The rest of the proof directly follows from the proof of Theorem 1 in [10]. ∎
@@
Let $\widetilde{H}_{2\text{-}D}$ be a parity-check tensor obtained by selecting first $w$ ($1\leq w\leq p$) horizontal block-layers from $H_{2\text{-}D}$. Then, there exists a 2-D classical QC-LDPC code with parameters $p^{4},(p^{2}-w+1)(p^{2}-1)]_{2}$ whose Tanner graph is free of cycles of length $4$. Moreover, the constructed code has the capability to correct $p\times p$ burst erasures.
@@
Theorem 2 establishes the dimension of the proposed code, while Lemma 4 guarantees that the Tanner graph of the code is free of cycles of length $4$.
@@
II-B 2-D Classical QC-LDPC Codes by Tiling $p\times p\times p$ Permutation Tensors ($p$ composite number)
In the previous subsection, we have only constructed 2-D classical LDPC codes by tiling tensors of size $p\times p\times p$, where $p$ is an odd prime. Now, we are going to state some results, where the size of the tensor is $p\times p\times p$, where $p$ is composite.
@@
Let $c$, $b$ and $h$ be positive integers, where $p\mid h$. Let the parity-check tensor $H_{2\text{-}D}$ be constructed by stacking $c\times b\times h$ permutation tensors of size $p\times p\times p$ along the $i$-, $j$- and $k$-directions, with shifts defined in (8). If
@@
It is well known that the presence of short cycles—particularly cycles of length $4$ and 66—in the Tanner graph of a classical code adversely affects decoding performance. Thus far, our focus has been on constructing codes free of $4$-cycles. This naturally motivates the development of classical codes that avoid both $4$- and $6$-cycles. Accordingly, we proceed to construct 2-D classical QC-LDPC codes that are free of cycles of lengths $4$ and $6$. To this end, we first state the generalized Behrend sequence from [14].
For any integers $d\geq 1$, $n\geq 2$, $k\leq nd^{2}$ and $\delta\geq 2$, define the set $\mathcal{B}(n,k,d,\delta)$ of all numbers of the form
@@
 $\text{ and }p\geq\left(\frac{h}{p}-1\right)\left(\text{max}\{b\in\mathcal{B}(n,k,d,\delta))\}\right).$ (21)
@@
Let $H_{2\text{-}D}$ be a parity-check tensor of dimension $c\times b\times h$ along the $i$-, $j$-, and $k$-directions, stacked with shifts specified in (22), and let $\mathcal{C}$ denote the associated 2-D classical QC-LDPC code. Then the Tanner graph of $\mathcal{C}$ contains no cycles of length $4$ or $6$. Moreover, the constructed code is capable of correcting any $p\times p$ burst erasure.
@@
First, we prove that the Tanner graph of the code does not have $4$-cycles. To, prove this, consider arbitrary horizontal block-layers $k_{0}$ and $k_{1}$, where $0\leq k_{0},k_{1}\leq(\frac{h}{p}-1)$. To check the existence of $4$-cycle, consider the following closed path of tensors indices:
@@
where $i_{s}\in\{0,1,2,\ldots,(c-1)\}$, $j_{t}\in\{0,1,2,\ldots,(b-1)\}$ for all $s,t\in\{0,1\}$. Then, by Theorem 1, the $4$-cycle along the closed path $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0});(i_{1},j_{1},k_{1}),(i_{0},j_{0},k_{1})$ exists if and only if the following equations are satisfied:
$\sum_{m=0}^{1}\Bigl(a(i_{m},j_{m},k_{m})-a(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{1}\Bigl(b(i_{m},j_{m},k_{m})-b(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
@@
$\alpha((\phi(i_{0})-\phi(i_{1}))+(\psi(j_{0})-\psi(j_{1})))=0.$ (23)
$\alpha((\phi(i_{0})-\phi(i_{1}))=0.$ (24)
Since $\alpha=\left(\left\lfloor\frac{k_{1}}{p}\right\rfloor-\left\lfloor\frac{k_{0}}{p}\right\rfloor\right)\neq 0$. Consider the equation (24), by Lemma 1 in [14], we have $\phi(i_{0})=\phi(i_{1})$ which implies that $i_{0}=i_{1}$. By putting $i_{0}=i_{1}$ in equation (23), we have
@@
Next, we proceed to prove that the Tanner graph of $H_{2\text{-}D}$ is free from $6$-cycles. To prove that consider arbitrary horizontal block-layer $k_{0}$, $k_{1}$ and $k_{2}$, where $0\leq k_{0},k_{1},k_{2}\leq(\frac{h}{p}-1)$. To check the existence of $6$-cycle, consider the closed path of tensors indices: $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0});$ $(i_{1},j_{1},k_{1}),(i_{2},j_{2},k_{1});$ $(i_{2},j_{2},k_{2}),$ $(i_{0},j_{0},k_{2}),$ where $i_{s}\in\{0,1,2,\ldots,(c-1)\}$ and $j_{t}\in\{0,1,2,\ldots,(b-1)\}$ for all $s,t\in\{0,1\}$.
Then, by Theorem 1, the $6$-cycle along the considered closed path exists if and only if the following equations are satisfied:
$\sum_{m=0}^{2}\Bigl(a(i_{m},j_{m},k_{m})-a(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{2}\Bigl(b(i_{m},j_{m},k_{m})-b(i_{m+1},j_{m+1},k_{m})\Bigr)$ $=0\pmod{p}$,
@@
where $\alpha=\left(\left\lfloor\frac{k_{1}}{p}\right\rfloor-\left\lfloor\frac{k_{0}}{p}\right\rfloor\right)$, $\beta=\left(\left\lfloor\frac{k_{2}}{p}\right\rfloor-\left\lfloor\frac{k_{1}}{p}\right\rfloor\right)$ and $\gamma=\left(\left\lfloor\frac{k_{2}}{p}\right\rfloor-\left\lfloor\frac{k_{0}}{p}\right\rfloor\right)$. It is easy to see that $\gamma=\alpha+\beta$. Now, consider the equation (26)
@@
In this section, we construct two classes of 2-D entanglement-assisted quantum LDPC codes by emplyoing proposed 2-D classical QC-LDPC codes. The first class is constructed from two distinct 2-D classical QC-LDPC codes, and the second class is obtained by employing a single 2-D classical Qc-LDPC code. Notably, the construction of the first class requires only a single ebit, and the unassisted portion of the overall Tanner graph is free of cycles of length $4$.
Now, consider the parity-check tensor $H_{2\text{-}D}$ constructed using the shifts defined in (8). We partition $H_{2\text{-}D}$ into two sub parity-check tensors, denoted by $H^{(1)}_{2\text{-}D}$ and $H^{(2)}_{2\text{-}D}$. The first sub parity-check tensor consists of first $w_{1}$ horizontal block-layers, where $1\leq w_{1}\leq p$, and the second consists of $w_{2}$ horizontal block-layers, where $1\leq w_{2}\leq p$, such that $2\leq w_{1}+w_{2}\leq p$. Moreover, $H^{(1)}_{2\text{-}D}$ and $H^{(2)}_{2\text{-}D}$ do not share any common horizontal block-layer. We then obtain the following theorems based on [15, Corollary 1].
@@
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be 2-D classical QC-LDPC codes with $H^{(1)}_{2D}$ and $H^{(2)}_{2D}$ as their parity check tensors. Then there exists a 2-D entanglement-assisted quantum LDPC codes with the parameters $p^{4},p^{4}-2p^{2}-(p^{2}-1)(w_{1}+w_{2}-2)+1;1]]_{2}.$ Moreover, its unassisted portion of overall tanner graph is of girth >4>4.
@@
Let $\mathcal{C}$ be a 2-D classical QC-LDPC code and $\widetilde{H}_{2\text{-}D}$ be its parity check tensor. Then there exists a 2-D entanglement-assisted quantum LDPC codes, with the following parameters $p^{4},(p^{2}-1)(p^{2}-w+1);p^{2}+(w-1)(p^{2}-1)]]_{2}$. In addition, by construction the Tanner graph of $\mathcal{C}$ is free from cyles of length $4$.
@@
In this section, we have thus far constructed EA quantum LDPC codes by employing the proposed 2-D classical QC-LDPC codes with the prime parameter $p$. In a similar manner, 2-D EA quantum codes can also be constructed from the other 2-D classical QC-LDPC codes developed earlier. However, in those cases, analytically determining the code rate and required number of entangled bits becomes difficult.
@@ IV Conclusions
First, we derive a general condition for the existence of $2g$-cycles in 2-D classical QC-LDPC codes. We then construct a family of 2-D classical QC-LDPC codes with 2-D burst-erasure correction capability and determine their exact code rates. In addition, we show that their Tanner graphs are free of cycles of length $4$. Several other families are also constructed with higher girth and the same erasure-correction capability. We employ the constructed 2-D classical QC-LDPC codes to develop two families of entanglement-assisted quantum LDPC codes. In the first family, two distinct 2-D classical QC-LDPC codes are used to ensure that the unassisted portion of the overall Tanner graph is free of $4$-cycles, while requiring only a single ebit. The second family is constructed using a single 2-D classical QC-LDPC code, where the underlying classical code itself is free of $4$-cycles.

### v2

@@ I Introduction
Let $H_{2\text{-D}}$ be the parity-check tensor constructed based on the shifts in (8), and let $\ell_{1}$ and $\ell_{2}$ be any two horizontal layers in $H_{2\text{-D}}$, then we obtain
@@
Viewing the layers of $H_{2\text{-D}}$ as matrices, motivates the following definition.
@@
The maximum number of linearly independent layers in a parity-check tensor $H_{2\text{-D}}$ is referred to as its rank.
Using this definition, we now state, in the following theorem, the rank of a parity-check tensor $H_{2\text{-D}}$ formed by selecting a subset of horizontal block-layers.
@@
Taking the shifts defined in (8) into account, it is straightforward to verify that the sum of all layers within any horizontal block-layer is the all-ones matrix, which implies a linear dependence among the layers in horizontal block-layers. Now, consider the $w$ horizontal block-layers of $\widetilde{H}_{2\text{-D}}$, and remove one horizontal layer from every horizontal block-layer except the first one. We delete the last horizontal layer from every horizontal block-layer except the first one and consider the following linear combination of all the remaining layers in the all horizontal-block layers:
@@
where $L_{i}^{(j)}$ stands for the $i$th layer in the $j$th horizontal block-layer and the coefficients are binary scalars. Consider the inner product defined in (11) with $\pmod{2}$. The remainder of the proof follows directly from Lemma 4 and the same line of reasoning employed in the proof of Theorem 1 in [10]. ∎
@@
Let $\widetilde{H}_{2\text{-D}}$ be a parity-check tensor obtained by selecting first $w$ ($1\leq w\leq p$) horizontal block-layers from $H_{2\text{-D}}$. Then there exists a 2-D classical QC-LDPC code with parameters $p^{4},(p^{2}-w+1)(p^{2}-1)]_{2}$ whose Tanner graph is free of cycles of length $4$. Moreover, the constructed code has the capability to correct $p\times p$ burst erasures.
@@
Theorem 2 establishes the dimension of the proposed code, while in Lemma 3 we have already proved that the Tanner graph of the code is free of cycles of length $4$.
@@
In the previous subsection, we have only constructed 2-D classical LDPC codes by tiling tensors of size $p\times p\times p$, where $p$ is an odd prime. Now, we are going to state a result, where the size of the tensor is $p\times p\times p$, where $p$ is composite.
@@
Let $c$, $b$ and $h$ be positive integers, where $p\mid h$. Let the parity-check tensor $H_{2\text{-D}}$ be constructed by stacking $c\times b\times h$ permutation tensors of size $p\times p\times p$ along the $i$-, $j$- and $k$-directions, respectively with shifts defined in (8). If
@@
In the next subsection, we focus on the construction of 2-D classical QC-LDPC codes with girth greater than $6$.
II-B 2-D Classical QC-LDPC Codes with Girth Greater than 66
It is well-known that the presence of short cycles—particularly cycles of length 44—in the Tanner graph of a classical code adversely affects decoding performance. Thus far, our focus has been on constructing codes free of $4$-cycles. This naturally motivates the development of classical codes that avoid both $4$- and $6$-cycles. Accordingly, we proceed to construct 2-D classical QC-LDPC codes that are free of cycles of lengths $4$ and $6$. To this end, we first state the generalized Behrend sequence from [14].
For any integers $d\geq 1$, $n\geq 2$, $k\leq nd^{2}$ and $\delta\geq 2$, define the set $\mathcal{B}(n,k,d,\delta)$ as follows:
@@
$p\geq\left(\frac{h}{p}-1\right)\left(\text{max}\{b\in\mathcal{B}(n,k,d,\delta))\}\right)+1.$ (21)
@@
Let $H_{2\text{-D}}$ be a parity-check tensor of dimension $c\times b\times h$ along the $i$-, $j$-, and $k$-directions, respectively stacked with shifts specified in (22), and let $\mathcal{C}$ denote the associated 2-D classical QC-LDPC code. Then the Tanner graph of $\mathcal{C}$ contains no cycles of length $4$ or $6$. Moreover, the constructed code is capable of correcting any $p\times p$ burst erasure.
@@
First, we prove that the Tanner graph of the code does not have $4$-cycles. To, prove this, consider arbitrary horizontal block-layers $k_{0}$ and $k_{1}$, where $0\leq k_{0},k_{1}\leq(\frac{h}{p}-1)$. To check the existence of $4$-cycle, consider the following arbitrary closed path of length $4$ of tensors indices:
@@
where $i_{s}\in\{0,1,2,\ldots,(c-1)\}$, $j_{t}\in\{0,1,2,\ldots,(b-1)\}$ for all $s,t\in\{0,1\}$. Then, by Theorem 1, a $4$-cycle exists along the closed path $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0}),(i_{1},j_{1},k_{1}),(i_{0},j_{0},k_{1})$ if and only if there exists a set $\{\bar{k}_{0},\bar{k}_{1}\}$ of $k$-indices, where each $\bar{k}_{i}$ is the $k$-coordinate of a tensor in the horizontal block layer $k_{i}$, satisfying the conditions required to connect the corresponding cycle segments, and the following equations are satisfied simultaneously:
$\sum_{m=0}^{1}\Bigl(a(i_{m},j_{m},\bar{k}_{m})-a(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{1}\Bigl(b(i_{m},j_{m},\bar{k}_{m})-b(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
@@
$\alpha((\phi(i_{0})-\phi(i_{1}))+(\psi(j_{0})-\psi(j_{1})))=0\pmod{p}$, (23)
$\alpha((\phi(i_{0})-\phi(i_{1}))=0\pmod{p}$, (24)
where $\alpha=\left(\left\lfloor\frac{\bar{k}_{1}}{p}\right\rfloor-\left\lfloor\frac{\bar{k}_{0}}{p}\right\rfloor\right)\neq 0\pmod{p}$ and less than or equal to $\left(\frac{h}{p}-1\right)$. Consider the equation (24), by Lemma 1 in [14], we have $\phi(i_{0})=\phi(i_{1})$ which implies that $i_{0}=i_{1}$. By putting $i_{0}=i_{1}$ in equation (23), we have
@@
Next, we proceed to prove that the Tanner graph of $H_{2\text{-D}}$ is free from $6$-cycles. To prove that consider arbitrary horizontal block-layers $k_{0}$, $k_{1}$ and $k_{2}$, where $0\leq k_{0},k_{1},k_{2}\leq(\frac{h}{p}-1)$. To check the existence of $6$-cycle, consider the following arbitrary closed path of length $6$ of tensors indices: $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0});$ $(i_{1},j_{1},k_{1}),(i_{2},j_{2},k_{1});$ $(i_{2},j_{2},k_{2}),$ $(i_{0},j_{0},k_{2}),$ where $i_{s}\in\{0,1,2,\ldots,(c-1)\}$ and $j_{t}\in\{0,1,2,\ldots,(b-1)\}$ for all $s,t\in\{0,1\}$.
Then, by Theorem 1, a $6$-cycle exists along the closed path $(i_{0},j_{0},k_{0}),(i_{1},j_{1},k_{0});$ $(i_{1},j_{1},k_{1}),(i_{2},j_{2},k_{1});$ $(i_{2},j_{2},k_{2}),$ $(i_{0},j_{0},k_{2}),$ if and only if there exists a set $\{\bar{k}_{0},\bar{k}_{1},\bar{k}_{2}\}$ of $k$-indices, where each $\bar{k}_{i}$ is the $k$-coordinate of a tensor in the horizontal block layer $k_{i}$, satisfying the conditions required to connect the corresponding cycle segments, and the following equations are simultaneously satisfied:
$\sum_{m=0}^{2}\Bigl(a(i_{m},j_{m},\bar{k}_{m})-a(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
$\sum_{m=0}^{2}\Bigl(b(i_{m},j_{m},\bar{k}_{m})-b(i_{m+1},j_{m+1},\bar{k}_{m})\Bigr)$ $=0\pmod{p}$,
@@
where $\alpha=\left(\left\lfloor\frac{\bar{k}_{1}}{p}\right\rfloor-\left\lfloor\frac{\bar{k}_{0}}{p}\right\rfloor\right)$, $\beta=\left(\left\lfloor\frac{\bar{k}_{2}}{p}\right\rfloor-\left\lfloor\frac{\bar{k}_{1}}{p}\right\rfloor\right)$ and $\gamma=\left(\left\lfloor\frac{\bar{k}_{2}}{p}\right\rfloor-\left\lfloor\frac{\bar{k}_{0}}{p}\right\rfloor\right)$. It is easy to see that $\gamma=\alpha+\beta$. Now, consider the equation (26)
@@
In this section, we construct two classes of 2-D entanglement-assisted quantum LDPC codes by emplyoing proposed 2-D classical QC-LDPC codes. The first class is constructed from two distinct 2-D classical QC-LDPC codes, and the second class is obtained by employing a single 2-D classical QC-LDPC code. Notably, the construction of the first class requires only a single ebit, and the unassisted portion of the overall Tanner graph is free of cycles of length $4$. To understand the meaning of the unassisted portion of the overall graph, the reader may refer to [10].
Now, consider the parity-check tensor $H_{2\text{-D}}$ of dimension $p\times p\times p^{2}$, where stacked permutation tensors are of size $p\times p\times p$, constructed using the shifts defined in (8). We partition $H_{2\text{-D}}$ into two sub parity-check tensors, denoted by $H^{(1)}_{2\text{-D}}$ and $H^{(2)}_{2\text{-D}}$. The first sub parity-check tensor consists of first $w_{1}$ horizontal block-layers, where $1\leq w_{1}\leq(p-1)$, and the second consists of $w_{2}$ horizontal block-layers, where $1\leq w_{2}\leq(p-1)$, such that $2\leq w_{1}+w_{2}\leq p$. Moreover, $H^{(1)}_{2\text{-D}}$ and $H^{(2)}_{2\text{-D}}$ do not share any common horizontal block-layer. We then obtain the following theorems based on [15, Corollary 1].
@@
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be 2-D classical QC-LDPC codes with $H^{(1)}_{2\text{-D}}$ and $H^{(2)}_{2\text{-D}}$ as their parity-check tensors. Then there exists a 2-D EA quantum LDPC code with the parameters $p^{4},p^{4}-2p^{2}-(p^{2}-1)(w_{1}+w_{2}-2)+1;1]]_{2}.$ Moreover, its unassisted portion of overall Tanner graph is of girth >4>4.
@@
Let $\mathcal{C}$ be a 2-D classical QC-LDPC code and $\widetilde{H}_{2\text{-D}}$ be its parity-check tensor having distinct $w$ number of horizontal block-layers. Then there exists a 2-D EA quantum LDPC code, with the parameters $p^{4},(p^{2}-1)(p^{2}-w+1);p^{2}+(w-1)(p^{2}-1)]]_{2}$. In addition, by construction the Tanner graph of $\mathcal{C}$ is free from cyles of length $4$.
@@
In this section, we have thus far constructed EA quantum LDPC codes by employing the proposed 2-D classical QC-LDPC codes with the prime parameter $p$. In a similar manner, 2-D EA quantum codes can also be constructed from the other 2-D classical QC-LDPC codes developed earlier. However, in those cases, analytically determining the code rate and the required number of entangled bits becomes difficult.
@@ IV Conclusions
First, we derive a general condition for the existence of $2g$-cycle in 2-D classical QC-LDPC code. We then construct a family of 2-D classical QC-LDPC codes with 2-D burst-erasure correction capability and determine their exact code rates. In addition, we show that their Tanner graphs are free of cycles of length $4$. Several other families are also constructed with higher girth and the same erasure-correction capability. We employ the constructed 2-D classical QC-LDPC codes to develop two families of entanglement-assisted quantum LDPC codes. In the first family, two distinct 2-D classical QC-LDPC codes are used to ensure that the unassisted portion of the overall Tanner graph is free of $4$-cycles, while requiring only a single ebit. The second family is constructed using a single 2-D classical QC-LDPC code, where the underlying classical code itself is free of $4$-cycles.